OpenStudy (anonymous):

simplify the expresson (2/9)^2 a. 4/81 b. 4/18 c. 4/11 d. 2/9

OpenStudy (aripotta):

\[\frac{ 2 }{ 9 }\times \frac{ 2 }{ 9 }=~?\]

OpenStudy (anonymous):

a?

OpenStudy (aripotta):

yep